OpenStudy (xxkesha):

What is the derivative of cuberoot of ax^2 + cuberoot of a^2x ? Can you also give me the value of x if x=a?

OpenStudy (3mar):

First of all, let's type in this equation.

OpenStudy (3mar):

\[f(x)=\sqrt[3]{ax^2}+\sqrt[3]{a^{2x}}\] Tell me that it is correct

OpenStudy (xxkesha):

No the x in second term is not an exponent

OpenStudy (3mar):

\[f(x)=\sqrt[3]{ax^2}+\sqrt[3]{a^2x}\] Looks right?

OpenStudy (xxkesha):